If sin-1 x + sin-1 y + sin-1 z =3π/2 then x^100+y^100+z^100-9/(x^101+y^101+z^101)= Options are. a)0 b)1 c)2 d)3

Grade:12

2 Answers

Arun
6 years ago
Dear student
 
Sin-1 x + sin-1 y + sin-1 z = 3\pi/2
Means
x = y = z = 1
Hence
Expression becomes
(1+1+1 -9)/(1+1+1) 
= -6/3
= -2
 
Regards
Arun (askIITians forum expert)
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Sin-1 x + sin-1 y + sin-1 z = 3\pi/2
Means
x = y = z = 1
Hence
Expression becomes
(1+1+1 -9)/(1+1+1) 
= -6/3
= -2
 
Thanks and regards,
Kushagra